# Learning to Reason as Action Abstractions with Scalable Mid-Training RL

### v1

@@ Abstract
[Correspondence]Shenao Zhang: shenao@u.northwestern.edu; John Peebles: john@johnpeebles.com; Zirui Wang: ziruiw@apple.com
@@ 3.1 Mid-Training Acquires Policy Priors via Action Space Pruning
Mid-training prunes the action space by assigning low probability to actions unlikely to be useful across tasks. Formally, mid-training algorithms aim to prune away all suboptimal actions, defined as follows.
Definition 3.3 (Suboptimal Action).
Let $\mathcal{Z}_{\epsilon}(\mathcal{M})$ denote the union of $\epsilon$-optimal action subsets in $\mathcal{M}$. An action $z\in\mathcal{Z}$ is $(\epsilon,\sigma)$-suboptimal if $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\mathmybb{1}(z\notin\mathcal{Z}_{\epsilon}(\mathcal{M}))]\geq\sigma$.
@@
$\overline{\mathcal{Z}}\subseteq\mathcal{Z}$ is an $\epsilon$-optimal action subset if it is $\epsilon$-optimal for all tasks. Let $|\overline{\mathcal{Z}}|$ denote the minimal size of such a subset.
Notably, $\overline{\mathcal{Z}}$ always exists and $|\overline{\mathcal{Z}}|$ is finite since $\Delta(\mathcal{M},\mathcal{A})=0$ and $\mathcal{A}$ itself is an optimal action set. Now we are ready to give our result on the pruning efficiency during mid-training.
@@
Denote $|\mathcal{D}_{E}|$ as the number of expert demonstrations during mid-training and $\widehat{\mathcal{Z}}$ as the resulting pruned action space. If
$|\mathcal{D}_{E}|=\Theta\Bigl(|\overline{\mathcal{Z}}|\log(|\mathcal{Z}|/\delta)/\sigma\Bigr),$
then with probability at least $1-\delta$, all the $(\epsilon,\sigma)$-suboptimal actions can be pruned away from $\mathcal{Z}$, and the pruning error in Lemma 3.2 satisfies $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\Delta(\mathcal{M},\widehat{\mathcal{Z}})]\leq\epsilon$.
Theorem 3.5 shows that the number of expert samples required to prune suboptimal actions decreases as both $|\overline{\mathcal{Z}}|$ and $|\mathcal{Z}|$ shrink. With a higher pruning efficiency, i.e., a smaller $|\overline{\mathcal{Z}}|$ and $|\mathcal{Z}|$, the probability that suboptimal actions survive mid-training decreases, leading to a smaller pruning error. This result highlights the importance of a compact decision space and motivates defining $\mathcal{Z}$ as a space of action abstractions instead of primitive actions $\mathcal{A}$.
@@
Compared to the space of primitive actions $\mathcal{A}$, the action abstraction has a substantially smaller near-optimal action subset $|\overline{\mathcal{Z}}|$, since each action abstraction corresponds to a transferable skill spanning multiple tasks. Theorem 3.5 implies that learning directly at the primitive action level demands significantly more expert data, while temporal abstractions enable mid-training to more efficiently approximate a near-optimal action set and provide a stronger policy prior, thereby reducing the burden on post-training RL.
@@ 3.2 Mid-Training Accelerates RL Convergence
The analysis above focuses on pruning error while holding the post-training RL error fixed. However, even with zero pruning error (e.g., by retaining the full primitive action space $\widehat{\mathcal{Z}}=\mathcal{A}$), RL may still be highly inefficient for long-horizon tasks due to the vast action space. We now analyze how mid-training influences the convergence of online RL. Our result is based on value iteration due to its simplicity.
@@
For any action subspace $\mathcal{Z}^{\prime}\subseteq\mathcal{Z}$, to achieve $\varepsilon$-optimality satisfying $\left\|V_{N}-V^{*}\right\|_{\infty}\leq\varepsilon$, the required number of iterations $N$ is lower-bounded by $N\geq\frac{1}{1-\overline{\gamma}}\log\frac{R_{\text{max}}}{\varepsilon(1-\gamma)}$, where $\overline{\gamma}=\sup_{s,0}\mathbb{E}[\gamma^{\tau}|s,z]\leq\gamma$ and $R_{\text{max}}=\max_{s,a}R(s,a)$.
The result reveals that reasoning structures acquired during mid-training affect convergence through the duration $\tau$ of the temporally-extended actions. Longer actions reduce $\overline{\gamma}$, leading to faster RL convergence than NTP-style mid-training, where $\tau=1$ and $\overline{\gamma}=\gamma$. This makes intuitive sense as each Bellman backup jumps across $\tau$ steps in one shot, which shortens the effective planning horizon and shrinks the error faster per iteration. Similar $1/(1-\overline{\gamma})$ dependency also appears in bounds for broader RL algorithms, such as policy gradient (Agarwal et al., 2021; Zhang et al., 2023), though we omit them due to additional assumptions.
@@ 4.2 Temporally-Consistent Latents as Action Abstractions
where $\alpha\in[0,1]$ is a hyperparameter, $\delta(\cdot)$ is the Dirac delta function, and $U(\cdot)$ is the uniform distribution over $\mathcal{Z}$. The delta mass helps preserve a temporally-consistent latent as a high-level action abstraction (reusing $z_{t-1}$), while the uniform term promotes diversity across latents.
@@ 5 RA3: A Scalable Mid-Training Algorithm
$\mathcal{D}_{\text{KL}}\Bigl(q(z_{t}|s_{t},z_{0:t-1})\,\,||\,\,p(z_{t}|s_{t},z_{0:t-1})\Bigr)=\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)-(1-q_{\text{act}})\mathcal{H}\Bigl(q(z_{t}|s_{t},z_{0:t-1})\Bigr),$
@@ Appendix A Proofs
Let $\mathcal{B}=\{\widehat{\mathcal{Z}}\subseteq\mathcal{Z}:|\widehat{\mathcal{Z}}|=|\overline{\mathcal{Z}}|\}$ be the set of all $|\overline{\mathcal{Z}}|$-length subsets. We care about the event that there exist some suboptimal subsets of actions that are not pruned away, i.e., survive:
$p\Biggl(\bigcup_{\mathcal{Z}^{\prime}\in\mathcal{B}}\text{survive}(\mathcal{Z}^{\prime})\Biggr)\leq\sum_{\mathcal{Z}^{\prime}\in\mathcal{B}}p\bigl(\text{survive}(\mathcal{Z}^{\prime})\bigr)\leq\binom{|\mathcal{Z}|}{|\overline{\mathcal{Z}}|}e^{-\sigma|\mathcal{D}_{E}|}=\Theta\Bigl(|\mathcal{Z}|^{|\overline{\mathcal{Z}}|}\Bigr)e^{-\sigma|\mathcal{D}_{E}|},$
@@
Plugging in $|\mathcal{D}_{E}|=\Theta(|\overline{\mathcal{Z}}|\log(|\mathcal{Z}|/\delta)/\sigma)$ gives us
@@
i.e., all the $(\epsilon,\sigma)$-sumoptimal action subsets $\mathcal{Z}^{\prime}$ are pruned away with probability at least $1-\delta$.
Then the result follows by rejecting all the actions that are not in the union of the subsets that survive. That is, with probability at least $1-\delta$, all the $(\epsilon,\sigma)$-suboptimal actions can be pruned away from $\mathcal{Z}$.
According to Definition A.1, any pruned action set $\widehat{\mathcal{Z}}$ satisfies $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\mathmybb{1}(\Delta(\mathcal{M},\widehat{\mathcal{Z}})\leq\epsilon)]\geq\sigma$. Thus, with probability at least $1-\delta$, the pruning error $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\Delta(\mathcal{M},\widehat{\mathcal{Z}})]\leq\epsilon$. ∎
@@
We first define the value iteration operator $\mathcal{T}$ in the temporally extended action space $\mathcal{Z}$ as
$(\mathcal{T}V)(s)=\max_{z\in\mathcal{Z}}\Bigl\{\mathbb{E}_{\tau,s^{\prime}}\bigl[R_{\tau}+\gamma^{\tau}V(s^{\prime})\mid s,z\bigr]\Bigr\},$
@@
By Banach’s fixed-point theorem, $\mathcal{T}$ has a unique fixed point $V^{*}$ and value iteration $V_{n+1}=\mathcal{T}V_{n}$ converges geometrically:
$\left\|V_{n}-V^{*}\right\|_{\infty}\leq\overline{\gamma}^{n}\left\|V_{0}-V^{*}\right\|_{\infty}\leq\overline{\gamma}^{n}R_{\text{max}}/(1-\gamma),$
where the last inequality holds since $V_{0}=0$ and the maximum value satisfies $V_{\max}=\sum_{t}\gamma^{t}R_{\text{max}}=R_{\text{max}}/(1-\gamma)$.
In order to get $\left\|V_{N}-V^{*}\right\|_{\infty}\leq\varepsilon$, we obtain
$N\geq\frac{1}{1-\overline{\gamma}}\log\frac{R_{\text{max}}}{\varepsilon(1-\gamma)}.$
@@
$\mathcal{D}_{\text{KL}}\Bigl(q(z_{t}|s_{t},z_{0:t-1})\,\,||\,\,p(z_{t}|s_{t},z_{0:t-1})\Bigr)$
$\qquad=q_{\text{act}}\log\frac{q_{\text{act}}}{\alpha}+\sum_{z_{t}\neq\langle\text{act}\rangle}(1-q_{\text{act}})\log\frac{q(z_{t}|s_{t},z_{0:t-1})}{p(z_{t}|s_{t},z_{0:t-1})}$
$\qquad=q_{\text{act}}\log\frac{q_{\text{act}}}{\alpha}+\sum_{z_{t}\neq\langle\text{act}\rangle}(1-q_{\text{act}})\log\frac{q(z_{t}\mid s_{t},z_{0:t-1})}{(1-\alpha)U(z_{t})}$
$\qquad=q_{\text{act}}\log\frac{q_{\text{act}}}{\alpha}+\sum_{z_{t}\neq\langle\text{act}\rangle}(1-q_{\text{act}})\biggl[\log\frac{q(z_{t}\mid s_{t},z_{0:t-1})}{1-q_{\text{act}}}+\log\frac{1-q_{\text{act}}}{1-\alpha}-\log U(z_{t})\biggr]$
$\qquad=\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)-(1-q_{\text{act}})\mathcal{H}\Bigl(q(z_{t}|s_{t},z_{0:t-1})\Bigr),$
where the terms that are constant w.r.t. $q$ is omitted. ∎

### v3

@@ Abstract
[Correspondence]Shenao Zhang shenao@u.northwestern.edu; John Peebles papers@johnpeebles.com; Zirui Wang ziruiw@apple.com
@@ 3.1 Mid-Training Acquires Policy Priors via Action Space Pruning
$\overline{\mathcal{Z}}_{\epsilon}\subseteq\mathcal{Z}$ is an $\epsilon$-optimal action subset if it is $\epsilon$-optimal for all tasks. Let $|\overline{\mathcal{Z}}_{\epsilon}|$ denote the minimal size of such a subset.
Notably, $|\overline{\mathcal{Z}}_{\epsilon}|$ is always finite for any $\epsilon\geq 0$ because even when $\epsilon=0$, we may still take $\overline{\mathcal{Z}}_{0}=\mathcal{Z}$ as $\Delta(\mathcal{M},\mathcal{Z})=0$ for all $\mathcal{M}$. In order words, in the worst case, no pruning occurs and all the actions in the finite $\mathcal{Z}$ are retained. We now present our result on pruning efficiency during mid-training.
@@
Assume the rewards are within the range of $[0,1]$. Denote $|\mathcal{D}_{E}|$ as the number of expert demonstrations during mid-training and $\widehat{\mathcal{Z}}$ as the resulting pruned action space. If
$|\mathcal{D}_{E}|=\Theta\Bigl(|\overline{\mathcal{Z}}_{\epsilon}|\log(|\mathcal{Z}|/\delta)/\sigma\Bigr),$
then with probability at least $1-\delta$, the pruning error in Lemma 3.2 satisfies $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\Delta(\mathcal{M},\widehat{\mathcal{Z}})]\leq\epsilon(1-\sigma)+\sigma/(1-\gamma)$.
Theorem 3.4 shows that the number of expert samples required to prune suboptimal actions decreases as both $|\overline{\mathcal{Z}}_{\epsilon}|$ and $|\mathcal{Z}|$ shrink. With a higher pruning efficiency, i.e., a smaller $|\overline{\mathcal{Z}}_{\epsilon}|$ and $|\mathcal{Z}|$, the probability that suboptimal actions survive mid-training decreases, leading to a smaller pruning error. This result highlights the importance of a compact decision space and motivates defining $\mathcal{Z}$ as a space of action abstractions instead of primitive actions $\mathcal{A}$.
@@
Compared to the space of primitive actions $\mathcal{A}$, the space of action abstraction has a substantially smaller size of $|\mathcal{Z}|$ and $|\overline{\mathcal{Z}}_{\epsilon}|$, since each action abstraction corresponds to a transferable skill spanning multiple tasks. Theorem 3.4 implies that learning directly at the primitive action level demands significantly more expert data, while temporal abstractions enable mid-training to more efficiently approximate a near-optimal action set and provide a stronger policy prior, thereby reducing the burden on post-training RL.
@@ 3.2 Mid-Training Accelerates RL Convergence
The analysis above focuses on pruning error while holding the post-training RL error fixed. However, even with zero pruning error (e.g., by retaining the full primitive action space $\widehat{\mathcal{Z}}=\mathcal{A}$), RL may still be highly inefficient for long-horizon tasks due to the vast action space. We now analyze how mid-training influences the convergence of online RL. Our result is based on policy iteration due to its simplicity, which iteratively updates the policy to optimality w.r.t. the policy value.
@@
After at most $N=\mathcal{O}((|\mathcal{S}||\widehat{\mathcal{Z}}|-|\mathcal{S}|)/(1-\overline{\gamma}))$ iterations, the policy achieves optimality in the pruned action space $\widehat{\mathcal{Z}}$, where $\overline{\gamma}=\sup_{s,z}\mathbb{E}[\gamma^{\tau}|s,z]\leq\gamma$.
The result reveals that reasoning structures acquired during mid-training affect convergence through both the pruned action space size and the duration $\tau$ of the temporally-extended actions. Longer actions reduce $\overline{\gamma}$, leading to faster RL convergence than NTP-style mid-training, where $\tau=1$ and $\overline{\gamma}=\gamma$. This makes intuitive sense as each Bellman backup jumps across $\tau$ steps in one shot, which shortens the effective planning horizon and shrinks the error faster per iteration. Similar dependency on the effective planning horizon also appears in bounds for broader RL algorithms, such as policy gradient (Agarwal et al., 2021; Zhang et al., 2023), though we omit them due to additional assumptions. Moreover, the number of suboptimal actions that RL needs to eliminate during online interaction reduces with a more compact action space $\widehat{\mathcal{Z}}$.
@@ 4.2 Temporally-Consistent Latents as Action Abstractions
where $\alpha\in[0,1]$ is a hyperparameter, $\delta(\cdot)$ is the Dirac delta function, and $U(\cdot)$ is the uniform distribution over the finite set $\mathcal{Z}$. The delta mass helps preserve a temporally-consistent latent as a high-level action abstraction (reusing $z_{t-1}$), while the uniform term promotes diversity across latents.
@@ 5 RA3: A Scalable Mid-Training Algorithm
$\mathcal{D}_{\text{KL}}\Bigl(q(z_{t}|s_{t},z_{0:t-1})\,\,||\,\,p(z_{t}|s_{t},z_{0:t-1})\Bigr)=\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)-\mathcal{H}\Bigl(q(z_{t}|s_{t},z_{0:t-1})\Bigr)+\text{const},$
@@ Appendix A Proofs
Let $\mathcal{B}=\{\widehat{\mathcal{Z}}\subseteq\mathcal{Z}:|\widehat{\mathcal{Z}}|=|\overline{\mathcal{Z}}_{\epsilon}|\}$ be the set of all $|\overline{\mathcal{Z}}_{\epsilon}|$-size subsets. We care about the event that there exist some suboptimal subsets of actions that are not pruned away, i.e., survive:
$p\Biggl(\bigcup_{\mathcal{Z}^{\prime}\in\mathcal{B}}\text{survive}(\mathcal{Z}^{\prime})\Biggr)\leq\sum_{\mathcal{Z}^{\prime}\in\mathcal{B}}p\bigl(\text{survive}(\mathcal{Z}^{\prime})\bigr)\leq\binom{|\mathcal{Z}|}{|\overline{\mathcal{Z}}_{\epsilon}|}e^{-\sigma|\mathcal{D}_{E}|}=\Theta\Bigl(|\mathcal{Z}|^{|\overline{\mathcal{Z}}_{\epsilon}|}\Bigr)e^{-\sigma|\mathcal{D}_{E}|},$
@@
Plugging in $|\mathcal{D}_{E}|=\Theta(|\overline{\mathcal{Z}}_{\epsilon}|\log(|\mathcal{Z}|/\delta)/\sigma)$ gives us
@@
i.e., all the $(\epsilon,\sigma)$-suboptimal action subsets $\mathcal{Z}^{\prime}$ are pruned away with probability at least $1-\delta$.
In other words, any action subset $\widehat{\mathcal{Z}}$ after pruning is not $(\epsilon,\sigma)$-suboptimal with high probability, i.e., it satisfies $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\mathmybb{1}(\Delta(\mathcal{M},\widehat{\mathcal{Z}})>\epsilon)]<\sigma$, or equivalently $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\mathmybb{1}(\Delta(\mathcal{M},\widehat{\mathcal{Z}})\leq\epsilon)]\geq 1-\sigma$, with probability at least $1-\delta$.
Since the rewards are in $[0,1]$, we know that $V_{\mathcal{M}}^{*}(s_{0})\leq\sum_{t=0}^{\infty}\gamma^{t}=1/(1-\gamma)$. Thus, $\Delta(\mathcal{M},\widehat{\mathcal{Z}})\leq 1/(1-\gamma)$.
Therefore, with probability at least $1-\delta$, the pruning error $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M})}[\Delta(\mathcal{M},\widehat{\mathcal{Z}})]\leq\epsilon(1-\sigma)+\sigma/(1-\gamma)$. ∎
@@
We first define the operator $\mathcal{T}$ and $\mathcal{T}^{\pi}$ in the $\mathcal{Z}$ action space as
$(\mathcal{T}V)(s)=\max_{z\in\mathcal{Z}}\Bigl\{\mathbb{E}_{\tau,s^{\prime}}\bigl[R_{\tau}+\gamma^{\tau}V(s^{\prime})\mid s,z\bigr]\Bigr\},\quad(\mathcal{T}^{\pi}V)(s)=\mathbb{E}_{z\sim\pi(\cdot\mid s)}\mathbb{E}_{\tau,s^{\prime}}\bigl[R_{\tau}+\gamma^{\tau}V(s^{\prime})\mid s,z\bigr],$
@@
In policy iteration, since $\pi_{n+1}$ is greedy w.r.t. $V^{\pi_{n}}$, we have that
$V^{\pi_{n}}=\mathcal{T}_{\pi_{n}}V^{\pi_{n}}\leq\mathcal{T}V^{\pi_{n}}=\mathcal{T}_{\pi_{n+1}}V^{\pi_{n}}\leq V^{\pi_{n+1}}.$
By induction, we have $\mathcal{T}^{n}V^{\pi_{0}}\leq V^{\pi_{n}}\leq V^{*}$. Taking component-wise absolute values and then the maximum over the states, we get that
$\left\|V^{*}-V^{\pi_{n}}\right\|_{\infty}\leq\left\|V^{*}-\mathcal{T}^{n}V^{\pi_{0}}\penalty 10000\ \right\|_{\infty}=\left\|\mathcal{T}^{n}V^{*}-\mathcal{T}^{n}V^{\pi_{0}}\penalty 10000\ \right\|_{\infty}\leq\overline{\gamma}^{n}\left\|V^{*}-V^{\pi_{0}}\right\|_{\infty}.$
Next, we show that after every $\mathcal{O}(1/(1-\overline{\gamma}))$ iterations, policy iteration eliminates at least one suboptimal state-action pair. For notation convenience, we define the operator $(M_{\pi}V)(s)=\mathbb{E}_{z\sim\pi(\cdot|s)}\mathbb{E}_{\tau,s^{\prime}}[\gamma^{\tau}V(s^{\prime})|s,z]$. Then we may write $\mathcal{T}_{\pi}V=r_{\pi}+M_{\pi}V$ and $V^{\pi}=(I-M_{\pi})^{-1}r_{\pi}$. Thus, we obtain that
$V^{*}-\mathcal{T}_{\pi_{n}}V^{*}=(I-M_{\pi_{n}})(V^{*}-V^{\pi_{n}})=(V^{*}-V^{\pi_{n}})-M_{\pi_{n}}(V^{*}-V^{\pi_{n}})\leq V^{*}-V^{\pi_{n}}.$
Taking the supremum gives us that
$\left\|V^{*}-\mathcal{T}_{\pi_{n}}V^{*}\right\|_{\infty}$ $\leq\left\|V^{*}-V^{\pi_{n}}\right\|_{\infty}\leq\overline{\gamma}^{n}\left\|V^{*}-V^{\pi_{0}}\right\|_{\infty}$
$=\overline{\gamma}^{n}\left\|(I-M_{\pi_{0}})^{-1}(V^{*}-\mathcal{T}_{\pi_{0}}V^{*})\right\|_{\infty}\leq\frac{\overline{\gamma}^{n}}{1-\overline{\gamma}}\left\|V^{*}-\mathcal{T}_{\pi_{0}}V^{*}\right\|_{\infty}.$
Let $\overline{s}\in\mathcal{S}$ be the state that satisfies $V^{*}(\overline{s})-\mathcal{T}_{\pi_{0}}V^{*}(\overline{s})=\left\|V^{*}-\mathcal{T}_{\pi_{0}}V^{*}\right\|_{\infty}$. Then we have
$V^{*}(\overline{s})-\mathcal{T}_{\pi_{n}}V^{*}(\overline{s})\leq\left\|V^{*}-\mathcal{T}_{\pi_{n}}V^{*}\right\|_{\infty}\leq\frac{\overline{\gamma}^{n}}{1-\overline{\gamma}}\Bigl(V^{*}(\overline{s})-\mathcal{T}_{\pi_{0}}V^{*}(\overline{s})\Bigr).$
Therefore, for some $n^{*}=\mathcal{O}(1/(1-\overline{\gamma}))$, when $n>n^{*}$, $\frac{\overline{\gamma}^{n}}{1-\overline{\gamma}}<1$. This indicates that
$\mathcal{T}_{\pi_{n}}V^{*}(\overline{s})>\mathcal{T}_{\pi_{0}}V^{*}(\overline{s}).$
Therefore, $\pi_{n}(\overline{s})\neq\pi_{0}(\overline{s})$, i.e., at least one suboptimal state-action pair is eliminated by the policy iteration algorithm. Since there are $|\mathcal{S}||\widehat{\mathcal{Z}}|-|\mathcal{S}|$ suboptimal state-action pairs in total, the policy achieves optimality after at most $\mathcal{O}((|\mathcal{S}||\widehat{\mathcal{Z}}|-|\mathcal{S}|)/(1-\overline{\gamma}))$ iterations. ∎
@@
$\mathcal{D}_{\text{KL}}\Bigl(q(z_{t}\mid s_{t},z_{0:t-1})\,\,||\,\,p(z_{t}\mid s_{t},z_{0:t-1})\Bigr)$ $=q_{\text{act}}\log\frac{q_{\text{act}}}{\alpha}+\sum_{z_{t}\neq\langle\text{act}\rangle}q(z_{t}\mid s_{t},z_{0:t-1})\log\frac{q(z_{t}\mid s_{t},z_{0:t-1})}{p(z_{t}\mid s_{t},z_{0:t-1})}$
$=q_{\text{act}}\log\frac{q_{\text{act}}}{\alpha}+\sum_{z_{t}\neq\langle\text{act}\rangle}q(z_{t}\mid s_{t},z_{0:t-1})\log\frac{q(z_{t}\mid s_{t},z_{0:t-1})}{(1-\alpha)U(z_{t})}.$ (A.4)
We define $\bar{q}(z_{t}|s_{t},z_{0:t-1})=q(z_{t}|s_{t},z_{0:t-1})/(1-q_{\text{act}})$ for $z_{t}\neq\langle\text{act}\rangle$. Then
$\sum_{z_{t}\neq\langle\text{act}\rangle}q(z_{t}\mid s_{t},z_{0:t-1})\log\frac{q(z_{t}\mid s_{t},z_{0:t-1})}{(1-\alpha)U(z_{t})}$
$\qquad=(1-q_{\text{act}})\Biggl[\sum_{z_{t}\neq\langle\text{act}\rangle}\bar{q}(z_{t}\mid s_{t},z_{0:t-1})\log\bar{q}(z_{t}\mid s_{t},z_{0:t-1})+\log\frac{1-q_{\text{act}}}{1-\alpha}-\mathbb{E}_{\bar{q}}\log U(z_{t})\Biggr].$ (A.5)
From the KL divergence between Bernoulli distributions, we obtain that
$\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)=q_{\text{act}}\log\frac{q_{\text{act}}}{\alpha}+(1-q_{\text{act}})\log\frac{1-q_{\text{act}}}{1-\alpha}.$
Besides, we have
$\sum_{z_{t}\neq\langle\text{act}\rangle}\bar{q}(z_{t}\mid s_{t},z_{0:t-1})\log\bar{q}(z_{t}\mid s_{t},z_{0:t-1})-\mathbb{E}_{\bar{q}}\log U(z_{t})=\mathcal{D}_{\text{KL}}\Bigl(\bar{q}(z_{t}\mid s_{t},z_{0:t-1})\,\,||\,\,U(z_{t})\Bigl).$
By plugging the above two KL divergence terms into (A.5) and (A.5), we obtain
$\mathcal{D}_{\text{KL}}\Bigl(q(z_{t}\mid s_{t},z_{0:t-1})\,\,||\,\,p(z_{t}\mid s_{t},z_{0:t-1})\Bigr)$
$\quad=\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)-(1-q_{\text{act}})\mathcal{D}_{\text{KL}}\Bigl(\bar{q}(z_{t}\mid s_{t},z_{0:t-1})\,\,||\,\,U(z_{t})\Bigl)$
$\quad=\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)-\sum_{z_{t}\neq\langle\text{act}\rangle}q(z_{t}\mid s_{t},z_{0:t-1})\log\frac{q(z_{t}\mid s_{t},z_{0:t-1})}{U(z_{t})(1-q_{\text{act}})}$
$\quad=\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)-\sum_{z_{t}\neq\langle\text{act}\rangle}q(z_{t}\mid s_{t},z_{0:t-1})\log q(z_{t}\mid s_{t},z_{0:t-1})+(1-q_{\text{act}})\log(1-q_{\text{act}})+\text{const}$
$\quad=\mathcal{D}_{\text{KL}}\Bigl(\text{Bern}(q_{\text{act}})\,\,||\,\,\text{Bern}(\alpha)\Bigl)-\mathcal{H}\Bigl(q(z_{t}\mid s_{t},z_{0:t-1})\Bigr)+\text{const},$
where the constant comes from the uniform prior term that appears when expanding the log and can be omitted. This gives us the desired result. ∎